how many liters of a face cream that cost $3360 per liter must be mixed with 12 liters of face cream that cost $1120 per liter to make a face cream that cost $1680 per liter?

Answer by longjonsilver(2297)

Let x = number of litres (English spelling) required.
(3360*x) + (1120*12) = 1680*(12+x)
3360x + 13440 = 20160 + 1680x
3360x - 1680x = 20160 - 13440
1680x = 6720
x = 6720/1680
x = 4litres

jon.